OpenStudy (anonymous):

a rectangular pool is 7 feet wide. it is 3 times long as it is wide. find the perimeter of the pool

OpenStudy (anonymous):

The Perimeter = P, Width = W Length = L P=2W+2L We know W=7 and L=3W or 3*7 if you plug those numbers in you get P=2*7=2(3*7) can you get it from there?